OpenStudy (anonymous):

If 2x + 5 = 9, what is the value of x?

OpenStudy (anonymous):

2x + 5 = 9 (subtract 5 from each side) 2x = 9 - 5 2x = 4 (now divide each side by 2) x = 2

OpenStudy (anonymous):

any questions ?